# Dr.E Bridges Graphs with Large Language Models through Words

### v2

@@ Abstract
Significant efforts have been dedicated to integrating the powerful Large Language Models (LLMs) with diverse modalities, particularly focusing on the fusion of language, vision and audio data. However, the graph-structured data, which is inherently rich in structural and domain-specific knowledge, has not yet been gracefully adapted to LLMs. Existing methods either describe the graph with raw text, suffering the loss of graph structural information, or feed Graph Neural Network (GNN) embeddings into LLMs at the cost of losing explainable prompt semantics. To bridge this gap, we introduce an end-to-end modality-aligning framework for LLM-graph alignment: Dual-Residual Vector Quantized-Variational AutoEncoder, namely Dr.E. Our approach is purposefully designed to facilitate token-level alignment with LLMs, enabling an effective translation of the intrinsic ‘language’ of graphs into comprehensible natural language. We also manage to enhance LLMs’ more robust structural understanding of graphs by incorporating multiple views of the central nodes based on their surrounding nodes at various distances. Our experimental evaluations on standard graph tasks demonstrate competitive performance against other state-of-the-art (SOTA) approaches. Additionally, our framework ensures certain visual interpretability, efficiency, and robustness, marking the promising successful endeavor to achieve token-level alignment between LLMs and GNNs. Our code is available at: https://anonymous.4open.science/r/dre-817.
@@ 1 Introduction
Graph data (Kipf and Welling 2016a), characterized by its domain-specific nature and diverse forms of features, is ubiquitous across various domains (Zhou et al. 2020; Wu et al. 2020, 2023b, 2023a; Zhao et al. 2023b, a). For instance, in citation networks (McCallum et al. 2000), it is used to predict the category of a given paper or the citation relationship between two papers. In recommendation systems (Zhao et al. 2023c), graph nodes represent items and users, and the system recommends items to users by mining the underlying relationships among them. Graph data usually lacks textual attributes, which poses challenges to constructing a uniform input representation for LLMs. Traditional graph-based methods (Spitzer 2001; Wu et al. 2020), have struggled to create a unified embedding space that aligns well with LLMs. Current methods (Ye et al. 2023; Wu et al. 2024) that integrate LLMs with graph data mostly employ one of two lines: it either directly leverages textual description of node information and graph structure, or it feeds the embeddings derived from GNNs into the LLMs.
Language serves as a natural medium to describe graphs, since LLMs are pretrained on massive textual datasets. Non-linguistic features often elude interpretation through verbal descriptions and, as a result, cannot be processed effectively by LLMs. An alternative strategy involves directly integrating graph embeddings into LLMs, but this approach does not fully exploit the linguistic interpretability of LLMs. Furthermore, within the graph convolution process, there is a risk of losing layer-specific information, which can be detrimental to the ability of LLMs to grasp layer-distinguishable structural knowledge effectively. Contrary to existing methods, we propose a novel solution that leverages the inherent linguistic processing capabilities of LLMs. By aligning graph embeddings with the LLMs’ vocabulary at different stages during the aggregation process, our approach enables LLMs to interpret graph structures linguistically. Given the intricacy of practical semantic information and the variability of graph structures, we have developed a codebook-driven method that employs multiple structural tokens (codes) to represent a single view of a graph’s structure. This approach effectively captures structural nuances through the iterative generation of new codes. In addition, considering the ever-changing graph structures, in order to obtain more stable structural tokens, we carefully deploy a structural encoder that focuses on multiple hop-different views, thereby achieving more robust graph-language model alignment and reducing training bias.
@@ 2.2 Intergration of LLM and GNN
It can broadly be categorized into two main approaches for integrating LLMs and GNNs (Chen et al. 2024b). The first category involves utilizing LLMs as enhancers to augment textual information, which is subsequently leveraged by GNNs for downstream tasks such as node classification. A representative example in this category is TAPE (He et al. 2023), which employs LLMs to generate pseudo-labels and explanatory content. These explanations elucidate the logical connections between textual features and their corresponding labels. Based on these explanations, textual features and enhanced textual features are derived from both the original and enriched text attributes, respectively, serving as the initial node features for the GNNs. SimTeG (Zhong and Chen 2020) on the other hand replaces the original embeddings from Pre-Trained Language Models (PLMs) with those obtained through Parameter-Efficient Fine-Tuning (PEFT), while Graphformers (Yang et al. 2021) promotes co-training between PLMs and GNNs through mutual generation of embeddings.
Instead of merely utilizing LLMs as annotators, it is more advantageous to engage them as predictors, where the final output of the model is presented entirely in natural language. An early attempt was to inject the graph embeddings directly into the prompt, thereby enabling LLMs to process node features, a technique employed by InstructGLM (Ye et al. 2023) through instruction-tuning of the LLMs. Later, GraphGPT (Tang et al. 2023) sought to add new tokens into LLMs vocabulary through a dual-stage tuning process. Methods such as LLaGA (Chen et al. 2024a), which avoided fine-tuning the LLMs, instead employ a projector to map graph embeddings to token embeddings, subtly encoding graph structural information. In contrast, the UniGraph model (He and Hooi 2024) not only trains a graph encoder but also fine-tunes an LLM to execute general GNN tasks using large language models.
@@ 3.2 Token-Level Alignment
where $\mathbf{e}_{c}$ represents the token embedding for index $c$, and $\mathcal{O}$ is the subset of the LLM’s vocabulary. Through this substitution, we achieve the transition from graph embeddings to language embeddings. During the fine-tuning phase of the large model, natural language is used as input. As illustrated in Table 6, we provide the potential node classification outcomes and the multi-view information of the central node, derived from the previous section, as prompts to the LLMs. Consequently, the model updates the parameters of both the GNN and the LLM synchronously, guided by the LLM’s loss function.
@@ 4.1 Experimental Settings
Table 1: The statistical information of the datasets is shown below, where ‘#Node’ and ‘#Edge’ represent the number of nodes and edges, respectively, and ‘Sparsity’ indicates the density of the graph.
Dataset #Node #Edge Sparsity
Cora 2,708 5,429 14.8065
Pubmed 19,717 44,338 2.2810
Arxiv 2,449,029 61,859,140 0.2063
@@
To enhance the reproducibility of our experiments and ensure a fair comparison with prior work, we utilize the Llama2-7B as our LLM decoder, which is also widely adopted in various other studies. However, this choice introduces biases when compared with other studies that use variants of LLaMA or other LLMs as foundational models, such as InstructGLM (Ye et al. 2023), GraphGPT (Tang et al. 2023), LLaGA (Touvron et al. 2023), and others. Our model incorporates three graph convolutional layers as an encoder to capture three views of representation when encoding a node into tokens. To prevent zero values from dominating the embeddings, we include an activation function and a quantizing module after dropout regularization. The convolution blocks are uniform and scalable, allowing them to be stacked multiple times. We employ cosine similarity to identify the nearest token, a method that has been shown to outperform a language model (LM) head layer in terms of convergence speed of the model. We implement LoRA PEFT adjustments for Llama2-7B and establish two distinct learning rates for the GNN encoder and LLM decoder, set at $1\times 10^{3}$ and $1\times 10^{-4}$, respectively, with a weight decay of $5\times 10^{-4}$. The hidden dimension for the SAGE convolution is 4096, matching the token embedding of LLaMA. Our experiments are conducted using 2 NVIDIA A800-SXM4-80GB GPUs.
@@
In the comparative experiments, we select three categories of methods. The first category encompasses classical approaches that are purely based on GNNs, including GCN (Kipf and Welling 2016a), GraphSage (Hamilton, Ying, and Leskovec 2017), GAT (Veličković et al. 2017), SGC (Wu et al. 2019), and SAGN (Sun, Gu, and Hu 2021). The second category comprises Transformer-based graph models, such as Nodeformer (Wu et al. 2022) and Graphformer (Wang et al. 2024). The third category includes hybrid models that integrate LLMs with GNNs, including GLEM (Zhao et al. 2022), InstructGLM (Ye et al. 2023), GraphMAE2 (Hou et al. 2023), Unigraph (He and Hooi 2024), GraphEdit (Guo et al. 2024), and LLaGA (Chen et al. 2024a). These experiments cover a sufficiently broad range to provide comprehensive insights into the comparative performance of these models.
@@ 4.3 Model Analysis
To evaluate the contribution of each component of Dr.E to the overall model performance, we design a series of ablation experiments, detailed in Table 3. These experiments consist of multiple iterations. Zero-Shot Prediction with Textual Information: We begin by extracting the textual information (titles and extractions) of nodes and performing zero-shot predictions using Llama2-7B. This demonstrates the LLM’s ability to comprehend textual attributes in graph data and classify papers based solely on their titles. Sequential Training of GNN and Frozen LLM: We then train the GNN encoder while keeping the LLM parameters frozen and updating only the GNN parameters. This sequential training improves the overall model performance, showing that a single GNN can align node embeddings with the input requirements of the language model to some degree. Fine-tuning LLM with LoRA: Building on this, we incorporate LoRA (Low-Rank Adaptation) for fine-tuning the LLM parameters, which leads to a significant improvement in performance. Multi-View Enhancement: Next, we apply the Multi-View Enhancement technique to structurally enhance the model’s performance, resulting in a significant improvement in scores. Quantization and Token Mapping: To integrate semantic information into the LLM, we quantize the embeddings and map them to tokens within Llama2-7B. However, this results in a slight decrease in performance, suggesting the need for additional adjustments. Dual-Layer Residual Mechanism and Codebook Refinement: Finally, we introduce the dual-layer residual mechanism and preprocess the codebook by filtering out tokens that do not carry meaningful semantic information. More details can be found in Appendix. With these improvements, the model achieves its best performance. These steps systematically evaluate the impact of each component on the overall effectiveness of Dr.E, highlighting critical aspects for optimizing the model’s performance.
@@
We also execute a case study, which involves analyzing the prediction for a real sample. In Table 6, we feed both the potential classification outcomes and the words represented by nodes into the LLM. The LLM then selects the most likely outcome based on the node information translated into natural language. This process reveals that while not every word carries positive label-related semantics, collectively, they contribute significantly to the final prediction. Although the words selected by our model do not necessarily convey practical semantic information, in some instances, Dr.E exhibits a certain degree of interpretability. As illustrated in Table 6, the terms ‘empirical’, ‘case’ and ‘statiscally’ are chosen to represent the graph structure for the central node, and the final classification for the node indeed reflects a ‘Case-Based’ approach. Additionly, other experiments and cases refer to Appendix.

### v3

# Multi-View Empowered Structural Graph Wordification for Language Models

@@ Abstract
Significant efforts have been dedicated to integrating the powerful Large Language Models (LLMs) with diverse modalities, particularly focusing on the fusion of language, vision and audio data. However, the graph-structured data, which is inherently rich in structural and domain-specific knowledge, has not yet been gracefully adapted to LLMs. Existing methods either describe the graph with raw text, suffering the loss of graph structural information, or feed Graph Neural Network (GNN) embeddings into LLMs at the cost of losing explainable prompt semantics. To bridge this gap, we introduce an end-to-end modality-aligning framework for LLM-graph alignment: Dual-Residual Vector Quantized-Variational AutoEncoder, namely Dr.E. Our approach is purposefully designed to facilitate token-level alignment with LLMs, enabling an effective translation of the intrinsic ‘language’ of graphs into comprehensible natural language. We also manage to enhance LLMs’ more robust structural understanding of graphs by incorporating multiple views of the central nodes based on their surrounding nodes at various distances. Our experimental evaluations on standard graph tasks demonstrate competitive performance against other state-of-the-art (SOTA) approaches. Additionally, our framework ensures certain visual interpretability, efficiency, and robustness, marking the promising successful endeavor to achieve token-level alignment between LLMs and GNNs. Our code is available at: https://github.com/Timothy914/Dr.E.
@@ 1 Introduction
Graph data (Kipf and Welling 2016a), characterized by its domain-specific nature and diverse forms of features, is ubiquitous across various domains (Zhou et al. 2020; Wu et al. 2020, 2023b, 2023a; Zhao et al. 2023b, a, 2024b). For instance, in citation networks (McCallum et al. 2000; Wu et al. 2024a), it is used to predict the category of a given paper or the citation relationship between two papers. In recommendation systems (Zhao et al. 2023c, 2024c, 2024a), graph nodes represent items and users, and the system recommends items to users by mining the underlying relationships among them. Graph data usually lacks textual attributes, which poses challenges to constructing a uniform input representation for LLMs. Traditional graph-based methods (Spitzer 2001; Wu et al. 2020), have struggled to create a unified embedding space that aligns well with LLMs. Current methods (Ye et al. 2023; Wu et al. 2024b) that integrate LLMs with graph data mostly employ one of two lines: it either directly leverages textual description of node information and graph structure, or it feeds the embeddings derived from GNNs into LLMs.
Language serves as a natural medium to describe graphs, since LLMs are pre-trained on massive textual datasets. Non-linguistic features often elude interpretation through verbal descriptions and, as a result, cannot be processed effectively by LLMs. An alternative strategy involves directly integrating graph embeddings into LLMs, but this approach does not fully exploit the linguistic interpretability of LLMs. Furthermore, within the graph convolution process, there is a risk of losing layer-specific information, which can be detrimental to the ability of LLMs to grasp layer-distinguishable structural knowledge effectively. Contrary to existing methods, we propose a novel solution that leverages the inherent linguistic processing capabilities of LLMs. By aligning graph embeddings with the LLMs’ vocabulary at different stages during the aggregation process, our approach enables LLMs to interpret graph structures linguistically. Given the intricacy of practical semantic information and the variability of graph structures, we have developed a codebook-driven method that employs multiple structural tokens (codes) to represent a single view of a graph’s structure. This approach effectively captures structural nuances through the iterative generation of new codes. In addition, considering the ever-changing graph structures, in order to obtain more stable structural tokens, we carefully deploy a structural encoder that focuses on multiple hop-different views, thereby achieving more robust graph-language model alignment and reducing training bias.
@@ 2.2 Intergration of LLM and GNN
It can be broadly classified into two main types of approaches for integrating LLMs and GNNs (Chen et al. 2024b). The first category involves utilizing LLMs as enhancers to augment textual information, which is subsequently leveraged by GNNs for downstream tasks such as node classification. A representative example in this category is TAPE (He et al. 2023), which employs LLMs to generate pseudo-labels and explanatory content. These explanations elucidate the logical connections between textual features and their corresponding labels. Based on these explanations, textual features and enhanced textual features are derived from both the original and enriched text attributes, respectively, serving as the initial node features for the GNNs. SimTeG (Duan et al. 2023) on the other hand replaces the original embeddings from Pre-Trained Language Models (PLMs) with those obtained through Parameter-Efficient Fine-Tuning (PEFT), while Graphformers (Yang et al. 2021) promotes co-training between PLMs and GNNs through mutual generation of embeddings.
Instead of merely utilizing LLMs as annotators, it is more advantageous to engage them as predictors, where the final output of the model is presented entirely in natural language. An early attempt was to inject the graph embeddings directly into the prompt, thereby enabling LLMs to process node features, a technique employed by InstructGLM (Ye et al. 2023) through instruction-tuning of the LLMs. Later, GraphGPT (Tang et al. 2024) sought to add new tokens into LLMs vocabulary through a dual-stage tuning process. Methods such as LLaGA (Chen et al. 2024a), which avoided fine-tuning the LLMs, instead employ a projector to map graph embeddings to token embeddings, subtly encoding graph structural information. In contrast, the UniGraph model (He and Hooi 2024) not only trains a graph encoder but also fine-tunes an LLM to execute general GNN tasks using large language models.
@@ 3.2 Token-Level Alignment
where $\mathbf{e}_{c}$ represents the token embedding for index $c$, and $\mathcal{O}$ is the subset of the LLM’s vocabulary. Through this substitution, we achieve the transition from graph embeddings to language embeddings. During the fine-tuning phase of the LLM, natural language is used as input. As illustrated in Table 4, we provide potential node classification outcomes and the multi-view information of the central node, derived from the previous section, as prompts to the LLMs. Consequently, the model updates the parameters of both the GNN and the LLM synchronously, guided by the LLM’s loss function.
@@ 4.1 Experimental Settings
Table 1: The statistical information of the datasets is shown below, where ‘#Node’ and ‘#Edge’ represent the number of nodes and edges, respectively, and ‘Sparsity‱’ indicates the density of the graph, computed as $S=1-\frac{|E|}{n(n-1)}$.
Dataset #Node #Edge Sparsity‱
Cora 2,708 5,429 14.8120
Pubmed 19,717 44,338 2.2811
Arxiv 2,449,029 61,859,140 0.2067
@@
To enhance the reproducibility of our experiments and ensure a fair comparison with prior work, we utilize the Llama2-7B as our LLM decoder, which is also widely adopted in various other studies. However, this choice introduces biases when compared with other studies that use variants of LLaMA or other LLMs as foundational models, such as InstructGLM (Ye et al. 2023), GraphGPT (tang2023graphgpt), LLaGA (Touvron et al. 2023), and others. Our model incorporates three graph convolutional layers as an encoder to capture three views of representation when encoding a node into tokens. To prevent zero values from dominating the embeddings, we include an activation function and a quantizing module after dropout regularization. The convolution blocks are uniform and scalable, allowing them to be stacked multiple times. We employ cosine similarity to identify the nearest token, a method that has been shown to outperform a language model (LM) head layer in terms of convergence speed of the model. We implement LoRA PEFT adjustments for Llama2-7B and establish two distinct learning rates for the GNN encoder and LLM decoder, set at $1\times 10^{-3}$ and $1\times 10^{-4}$, respectively, with a weight decay of $5\times 10^{-4}$. The hidden dimension for the SAGE convolution is 4096, matching the token embedding of Llama. Our experiments are conducted using 2 NVIDIA A800-SXM4-80GB GPUs.
@@
In the comparative experiments, we select three categories of methods. The first category encompasses classical approaches that are purely based on GNNs, including GCN (Kipf and Welling 2016a), GraphSage (Hamilton, Ying, and Leskovec 2017), GAT (Veličković et al. 2017), SGC (Wu et al. 2019), and SAGN (sun2021scalable). The second category comprises Transformer-based graph models, such as Nodeformer (Wu et al. 2022) and Graphformer (Wang et al. 2024). The third category includes hybrid models that integrate LLMs with GNNs, including GLEM (Zhao et al. 2022), InstructGLM (Ye et al. 2023), GraphMAE2 (Hou et al. 2023), Unigraph (He and Hooi 2024), GraphEdit (Guo et al. 2024), and LLaGA (Chen et al. 2024a). These experiments cover a sufficiently broad range to provide comprehensive insights into the comparative performance of these models.
@@ 4.3 Model Analysis
To evaluate the contribution of each component of Dr.E to the overall model performance, we design a series of ablation experiments, detailed in Table 3. These experiments consist of multiple iterations. Zero-Shot Prediction with Textual Information: We begin by extracting the textual information (titles and extractions) of nodes and performing zero-shot predictions using Llama2-7B. This demonstrates the LLM’s ability to comprehend textual attributes in graph data and classify papers based solely on their titles. Sequential Training of GNN and Frozen LLM: We then train the GNN encoder while keeping the LLMs’ parameters frozen and updating only the GNN parameters. This sequential training improves the overall model performance, showing that a single GNN can align node embeddings with the input requirements of the language model to some degree. Fine-tuning LLM with LoRA: Building on this, we incorporate Low-Rank Adaptation(LoRA) for fine-tuning the LLM, which leads to a significant improvement in performance. Multi-View Enhancement: Next, we apply the Multi-View Enhancement technique to structurally enhance the model’s performance, resulting in a significant improvement in scores. Quantization and Token Mapping: To integrate semantic information into the LLM, we quantize the embeddings and map them to tokens within Llama2-7B. However, this results in a slight decrease in performance, suggesting the need for additional adjustments. Dual-Layer Residual Mechanism and Codebook Refinement: Finally, we introduce the dual-layer residual mechanism and preprocess the codebook by filtering out tokens that do not carry meaningful semantic information. With these improvements, the model achieves its best performance. These steps systematically evaluate the influence of each component on the overall effectiveness of Dr.E, highlighting critical aspects for optimizing the model’s performance.
@@
We also execute a case study, which involves analyzing the prediction for a real sample. In Table 4, we feed both the potential classification outcomes and the words represented by nodes into the LLM. The LLM then selects the most likely outcome based on the node information translated into natural language. This process reveals that while not every word carries positive label-related semantics, collectively, they contribute significantly to the final prediction. Although the words selected by our model do not necessarily convey practical semantic information, in some instances, Dr.E exhibits a certain degree of interpretability. As illustrated in Table 4, the terms ‘empirical’, ‘case’ and ‘statistically’ are chosen to represent the graph structure for the central node, and the final classification for the node indeed reflects a ‘Case-Based’ approach.
@@ 5 Conclusion
6 Acknowledgments
This study was partially funded by the supports of National Natural Science Foundation of China (72471165, 72231004, 72101176).